# Cost-aware LLM-based Online Dataset Annotation

### v1

@@ Appendix C The Baseline Algorithm
4:     $\hat{y}_{t}\leftarrow\arg\max_{m\in[M]}\sum_{i=1}^{K}\omega_{\text{def},i}(t-1)\cdot{\mathds{1}\left\{y_{i,t}=m\right\}}$
5:     Generate rewards for LLMs: $r_{i,t}={\mathds{1}\left\{y_{i,t}=y_{t}\right\}}$
@@ Appendix D Supplementary Details for Experiments on the MMLU Dataset
Figure 3: Cumulative average accuracy (blue) and cost (red) of CaMVo ($k_{\min}=1$) on the MMLU dataset across rounds for various $\delta$. The green line marks each $\delta$-specific target accuracy.
D.1 Additional Experimental Results
Figure 3 illustrates CaMVo’s cumulative average accuracy (blue) and cost (red) over rounds for $k_{\min}=1$ under different confidence thresholds $\delta$ to explore CaMVo’s learning dynamics for various $\delta$ values. The green line marks each $\delta$-specific target accuracy. In all cases, the algorithm begins by querying larger, more expensive ensembles to gather reliable performance estimates, then swiftly transitions to cheaper subsets once the lower-confidence bounds stabilize. This yields a steep decline in cost concurrent with accuracy settling at a value above the target line. A temporary dip in accuracy around round 1,000 appears consistently, reflecting a cluster of harder instances in our fixed data shuffle.
For high thresholds ($\delta=0.99$), CaMVo predominantly queries the full ensemble, producing an almost linear cost profile. At intermediate levels ($\delta=0.98,0.975$), cost initially falls but momentarily rises when accuracy dips below the target, prompting the algorithm to select slightly costlier subsets to regain the required confidence as the accuracy estimations of individual LLMs decrease. When $\delta<0.965$, the cost curve decreases monotonically and converges to a stable minimum, indicating rapid identification of the context-specific optimal subsets.
Finally, for low thresholds ($\delta=0.85,0.80$), observed accuracy significantly exceeds the target owing to the performance gaps among individual LLMs: no model has true accuracy between 70% and 80%, hence CaMVo’s conservative lower-bound estimates result in consistently higher realized accuracy. Overall, these results underscore CaMVo’s capacity to balance exploration and exploitation, quickly pinpoint cost-effective ensembles, and reliably meet user-specified accuracy requirements.
Similarly, Figure 5 examines CaMVo’s learning curves for various $\delta$ values with $k_{\min}=3$. As before, CaMVo starts by querying larger, costlier ensembles to estimate model performance, then quickly shifts to cheaper subsets once the lower-confidence bounds converge. For $\delta\geq 0.95$, the trajectories closely mirror those with $k_{\min}=1$. When $\delta\leq 0.90$, however, both accuracy and cost stabilize even more rapidly—reflecting convergence to the least-expensive three-LLM ensemble. This demonstrates that increasing $k_{\min}$ enhances stability by preventing undersized subsets from being chosen at the expense of a modest cost increase.
@@ D.1 Additional Experimental Results
Refer to caption Refer to caption Refer to caption
Figure 4: Cumulative average accuracy (blue) and cost (red) of CaMVo ($k_{\min}=3$) on the MMLU dataset across rounds for various $\delta$. The green line marks each $\delta$-specific target accuracy.
Figure 5: Mean (solid lines) and one-standard-deviation bands (shading) of CaMVo’s cumulative average accuracy (blue) and cost (red) over 20 random shuffles of MMLU when $\delta=0.96$ , $k_{\min}=1$ (Left); and $\delta=0.96$ , $k_{\min}=1$ (Right). The green line indicates the accuracy target of $84.65\%$ for $\delta=0.96$.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 6: Cumulative average accuracy (blue) and cost (red) of CaMVo with $\delta=0.96$, $k_{\min}=1$ under nine different permutations of the dataset. The green line marks each $\delta$-specific target accuracy.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 7: Cumulative average accuracy (blue) and cost (red) of CaMVo with $\delta=0.96$, $k_{\min}=3$ under nine different permutations of the dataset. The green line marks each $\delta$-specific target accuracy.
Further, to evaluate CaMVo’s robustness to input ordering, Figure 5 shows the mean cumulative average accuracy and cost trajectories (solid lines) averaged over 20 random shuffles of the MMLU dataset for $\delta=0.96$, $k_{\min}=1$ (Left); and for $\delta=0.96$, $k_{\min}=3$ (Right). Shaded bands denote one standard deviation. Although the accuracy band is initially wide due to the exploration of CaMVo, and also different mixes of easy and hard examples across the shuffles; it contracts rapidly, underscoring CaMVo’s consistent attainment of the target accuracy across permutations. The cost band also narrows over time, illustrating stable convergence to low-cost ensembles. Notably, the accuracy band remains much tighter than the cost band, since CaMVo targets above the accuracy threshold but does not optimize for a fixed cost. Comparing $k_{\min}=1$ and $k_{\min}=3$, it can be seen that both the cost and accuracy bands are similar.
To evaluate CaMVo’s robustness to input ordering in more detail, Figure 7 overlays the mean cumulative average accuracy and cost plots of nine individual runs from these permutations. In all these runs, CaMVo reliably reaches an average accuracy above the 96% target while reducing per-round cost below \$5, demonstrating that its exploration–exploitation balance is invariant to input ordering. Figure 7 presents the corresponding results for the $k_{\min}=3$ setting, confirming similar robustness to input permutations.
@@ E.1 Additional Experimental Results
Figure 9: Cumulative average accuracy (blue) and cost (red) of CaMVo ($k_{\min}=3$) on the IMDB Movie Reviews Dataset across rounds for various confidence thresholds $\delta$. The green line marks each $\delta$-specific target accuracy.
Figure 10: Mean (solid lines) and one-standard-deviation bands (shading) of CaMVo’s cumulative average accuracy (blue) and cost (red) over 20 random shuffles of MMLU when $\delta=0.995$ , $k_{\min}=1$ (Left); and $\delta=0.995$ , $k_{\min}=1$ (Right). The green line indicates the accuracy target of $95.14\%$ for $\delta=0.995$.

### v2

@@ Appendix C The Baseline Algorithm
4:  $\hat{y}_{t}\leftarrow\arg\max_{m\in[M]}\sum_{i=1}^{K}\omega_{\text{def},i}(t)\cdot{\mathds{1}\left\{y_{i,t}=m\right\}}$
5:  Generate rewards for LLMs: $r_{i,t}={\mathds{1}\left\{y_{i,t}=\hat{y}_{t}\right\}}$
@@ Appendix D Supplementary Details for Experiments on the MMLU Dataset
D.1 Additional Experimental Results
$\bf\delta$ $\bf label$
Precision,
$\bf k_{\min}=1$
Recall,
$\bf k_{\min}=1$
F1 Score,
$\bf k_{\min}=1$
Precision,
$\bf k_{\min}=3$
Recall,
$\bf k_{\min}=3$
F1 Score,
$\bf k_{\min}=3$
0.99 0 0.87 0.86 0.87 0.87 0.86 0.87
0.99 1 0.87 0.89 0.88 0.87 0.89 0.88
0.99 2 0.88 0.9 0.89 0.88 0.9 0.89
0.99 3 0.91 0.89 0.9 0.91 0.89 0.9
0.98 0 0.88 0.86 0.87 0.88 0.86 0.87
0.98 1 0.88 0.89 0.88 0.88 0.89 0.88
0.98 2 0.88 0.9 0.89 0.88 0.9 0.89
0.98 3 0.91 0.9 0.9 0.91 0.9 0.9
0.975 0 0.88 0.85 0.87 0.88 0.85 0.87
0.975 1 0.87 0.89 0.88 0.87 0.89 0.88
0.975 2 0.88 0.9 0.89 0.88 0.9 0.89
0.975 3 0.91 0.9 0.9 0.91 0.9 0.9
0.97 0 0.88 0.85 0.86 0.88 0.85 0.86
0.97 1 0.87 0.89 0.88 0.87 0.89 0.88
0.97 2 0.88 0.9 0.89 0.88 0.9 0.89
0.97 3 0.91 0.9 0.9 0.91 0.9 0.9
0.965 0 0.89 0.85 0.87 0.89 0.85 0.87
0.965 1 0.86 0.88 0.87 0.86 0.88 0.87
0.965 2 0.88 0.9 0.89 0.88 0.9 0.89
0.965 3 0.91 0.9 0.9 0.91 0.9 0.9
0.96 0 0.88 0.85 0.86 0.88 0.85 0.86
0.96 1 0.86 0.88 0.87 0.86 0.88 0.87
0.96 2 0.87 0.89 0.88 0.87 0.89 0.88
0.96 3 0.9 0.89 0.9 0.9 0.89 0.9
0.955 0 0.87 0.84 0.86 0.87 0.84 0.86
0.955 1 0.86 0.88 0.87 0.86 0.88 0.87
0.955 2 0.87 0.88 0.88 0.87 0.88 0.88
0.955 3 0.89 0.89 0.89 0.89 0.89 0.89
0.95 0 0.87 0.84 0.85 0.87 0.84 0.85
0.95 1 0.85 0.87 0.86 0.85 0.87 0.86
0.95 2 0.87 0.87 0.87 0.87 0.87 0.87
0.95 3 0.89 0.89 0.89 0.89 0.89 0.89
0.9 0 0.85 0.82 0.84 0.85 0.82 0.84
0.9 1 0.84 0.85 0.84 0.84 0.85 0.84
0.9 2 0.85 0.85 0.85 0.85 0.85 0.85
0.9 3 0.86 0.87 0.87 0.86 0.87 0.87
0.85 0 0.84 0.82 0.83 0.84 0.82 0.83
0.85 1 0.81 0.86 0.84 0.81 0.86 0.84
0.85 2 0.85 0.84 0.84 0.85 0.84 0.84
0.85 3 0.87 0.85 0.86 0.87 0.85 0.86
0.8 0 0.8 0.82 0.81 0.8 0.82 0.81
0.8 1 0.76 0.87 0.81 0.76 0.87 0.81
0.8 2 0.85 0.8 0.82 0.85 0.8 0.82
0.8 3 0.88 0.8 0.84 0.88 0.8 0.84
0.75 0 0.59 0.73 0.65 0.59 0.73 0.65
0.75 1 0.71 0.65 0.68 0.71 0.65 0.68
0.75 2 0.76 0.63 0.69 0.76 0.63 0.69
0.75 3 0.72 0.74 0.73 0.72 0.74 0.73
0.7 0 0.58 0.73 0.65 0.58 0.73 0.65
0.7 1 0.71 0.64 0.67 0.71 0.64 0.67
0.7 2 0.75 0.63 0.68 0.75 0.63 0.68
0.7 3 0.71 0.74 0.73 0.71 0.74 0.73
Table 5: Precision, recall, and F1 scores of CaMVo with $k_{\min}=1$ and $k_{\min}=3$ across different confidence thresholds $\delta$ for each label category on the MMLU dataset.
$\bf label$ Precision Recall F1 Score
Maj. voting 0 0.87 0.86 0.86
Maj. voting 1 0.87 0.88 0.88
Maj. voting 2 0.88 0.89 0.89
Maj. voting 3 0.91 0.89 0.9
Table 6: Precision, recall and F1 score of majority voting over each output category on the MMLU dataset.
To facilitate a more detailed analysis of the results in Table 3, for each label category in Table 5 we report precision, recall, and F1 score as additional performance metrics. For comparison, the corresponding results for majority voting are presented in Table 6. These metrics allow to evaluate category-wise performance. The results indicate that category 0 is slightly more challenging to label, as reflected by its consistently lower scores across all metrics. Furthermore, as the confidence threshold $\delta$ decreases, all three metrics degrade across label categories, with a more pronounced decline observed for category 0, likely due to its possibly higher intrinsic difficulty.
@@ D.1 Additional Experimental Results
Figure 3: Cumulative average accuracy (blue) and cost (red) of CaMVo ($k_{\min}=1$) on the MMLU dataset across rounds for various confidence thresholds $\delta$. The green line marks each $\delta$-specific target accuracy.
Figure 3 illustrates CaMVo’s cumulative average accuracy (blue) and cost (red) over rounds for $k_{\min}=1$ under different confidence thresholds $\delta$ to explore CaMVo’s learning dynamics for various $\delta$ values. The green line marks each $\delta$-specific target accuracy. In all cases, the algorithm begins by querying larger, more expensive ensembles to gather reliable performance estimates, then swiftly transitions to cheaper subsets once the lower‐confidence bounds stabilize. This yields a steep decline in cost concurrent with accuracy settling at a value above the target line. A temporary dip in accuracy around round 1,000 appears consistently, reflecting a cluster of harder instances in our fixed data shuffle.
For high thresholds ($\delta=0.99$), CaMVo predominantly queries the full ensemble, producing an almost linear cost profile. At intermediate levels ($\delta=0.98,0.975$), cost initially falls but momentarily rises when accuracy dips below the target, prompting the algorithm to select slightly costlier subsets to regain the required confidence as the accuracy estimations of individual LLMs decrease. When $\delta<0.965$, the cost curve decreases monotonically and converges to a stable minimum, indicating rapid identification of the context‐specific optimal subsets.
Finally, for low thresholds ($\delta=0.85,0.80$), observed accuracy significantly exceeds the target owing to the performance gaps among individual LLMs: no model has true accuracy between 70% and 80%, hence CaMVo’s conservative lower‐bound estimates result in consistently higher realized accuracy. Overall, these results underscore CaMVo’s capacity to balance exploration and exploitation, quickly pinpoint cost‐effective ensembles, and reliably meet user‐specified accuracy requirements.
To evaluate CaMVo’s robustness to input ordering, Figure 4 shows the mean cumulative average accuracy and cost trajectories (solid lines) for $\delta=0.96$, $k_{\min}=1$, averaged over 20 random shuffles of the MMLU dataset. Shaded bands denote one standard deviation. Although the accuracy band is initially wide due to the exploration of CaMVo, and also different mixes of easy and hard examples across the shuffles; it contracts rapidly, underscoring CaMVo’s consistent attainment of the target accuracy across permutations. The cost band also narrows over time, illustrating stable convergence to low‐cost ensembles. Notably, the accuracy band remains much tighter than the cost band, since CaMVo targets above the accuracy threshold but does not optimize for a fixed cost.
Figure 4: Mean (solid lines) and one‐standard‐deviation bands (shading) of CaMVo’s cumulative average accuracy (blue) and cost (red) over 20 random shuffles of MMLU ($\delta=0.96$, $k_{\min}=1$). The green line indicates the accuracy target of $84.65\%$ for $\delta=0.96$.
@@
Figure 5: Cumulative average accuracy (blue) and cost (red) of CaMVo with $\delta=0.96$, $k_{\min}=1$ under nine different random permutations of the dataset. The green line marks each $\delta$-specific target accuracy.
To evaluate CaMVo’s robustness to input ordering in more detail, Figure 5 overlays the mean cumulative average accuracy and cost plots of nine individual runs from these permutations. In all these runs, CaMVo reliably reaches an average accuracy above the 96% target while reducing per‐round cost below \$5, demonstrating that its exploration–exploitation balance is invariant to input ordering.
@@ E.1 Additional Experimental Results
$\bf\delta$ $\bf label$
Precision
$\bf(k_{\min}=1)$
Recall
$\bf(k_{\min}=1)$
F1 Score
$\bf(k_{\min}=1)$
Precision
$\bf(k_{\min}=3)$
Recall
$\bf(k_{\min}=3)$
F1 Score
$\bf(k_{\min}=3)$
0.999 0 0.96 0.96 0.96 0.96 0.96 0.96
0.999 1 0.96 0.96 0.96 0.96 0.96 0.96
0.998 0 0.95 0.95 0.95 0.95 0.95 0.95
0.998 1 0.95 0.95 0.95 0.95 0.95 0.95
0.997 0 0.96 0.95 0.95 0.96 0.95 0.95
0.997 1 0.95 0.96 0.95 0.95 0.96 0.95
0.995 0 0.96 0.95 0.95 0.96 0.95 0.95
0.995 1 0.95 0.96 0.95 0.95 0.96 0.95
0.99 0 0.95 0.95 0.95 0.95 0.95 0.95
0.99 1 0.95 0.95 0.95 0.95 0.95 0.95
0.985 0 0.94 0.95 0.95 0.94 0.95 0.95
0.985 1 0.95 0.94 0.95 0.95 0.94 0.95
0.98 0 0.94 0.95 0.95 0.94 0.95 0.95
0.98 1 0.95 0.94 0.95 0.95 0.94 0.95
0.97 0 0.94 0.95 0.95 0.94 0.95 0.95
0.97 1 0.95 0.94 0.95 0.95 0.94 0.95
0.96 0 0.94 0.94 0.94 0.94 0.94 0.94
0.96 1 0.94 0.94 0.94 0.94 0.94 0.94
0.95 0 0.94 0.95 0.94 0.94 0.95 0.94
0.95 1 0.95 0.94 0.94 0.95 0.94 0.94
0.9 0 0.94 0.94 0.94 0.94 0.94 0.94
0.9 1 0.94 0.94 0.94 0.94 0.94 0.94
MJV 0 0.96 0.96 0.96
MJV 1 0.96 0.96 0.96
Table 7: Precision, recall, and F1 scores of CaMVo with $k_{\min}=1$ and $k_{\min}=3$ across different confidence thresholds $\delta$ for each output label on the IMDB dataset. Note that the last two rows report the corresponding precision, recall, and F1 scores obtained with majority voting.
To facilitate a more detailed analysis of the experimental results in Table 4, we report additional performance metrics—namely, precision, recall, and F1 score—for each label category in Table 7. These metrics allow for a finer-grained evaluation of CaMVo’s behavior across different output labels. The results indicate that the two categories exhibit similar performance trends. As expected, decreasing the confidence threshold $\delta$ leads to consistent degradation in all three metrics across both label categories.
@@
Figure 6 visualizes CaMVo’s learning trajectories on IMDB for $k_{\min}=1$ across all the confidence thresholds $\delta$ values reported in Table 4. In all cases, CaMVo begins by querying larger, higher‐cost ensembles to obtain reliable performance estimates, then rapidly shifts to cost-optimal subsets once the lower‐confidence bounds converge. This transition yields a sharp decline in cost while maintaining accuracy above the target line.
Figure 7: Mean (solid lines) and one‐standard‐deviation bands (shading) of CaMVo’s cumulative average accuracy (blue) and cost (red) over 20 random shuffles of MMLU ($\delta=0.995$, $k_{\min}=1$). The green line indicates the accuracy target of $95.14\%$ for $\delta=0.995$.
@@
Figure 8: Cumulative average accuracy (blue) and cost (red) of CaMVo with $\delta=0.96$, $k_{\min}=1$ under nine different random permutations of the dataset. The green line marks each $\delta$-specific target accuracy.
At the extreme threshold $\delta=0.999$, CaMVo predominantly queries the full ensemble, resulting in a near‐linear cost profile until about round 25,000. For $\delta\leq 0.98$, cost quickly converges to a stable minimum, reflecting identification of the least‐expensive subset that meets the target accuracy. Across all plots, CaMVo achieves or exceeds the respective accuracy target. For very high thresholds ($\delta\geq 0.995$), final accuracy hovers just above the threshold, as expected; as $\delta$ decreases, the accuracy surplus grows. Below $\delta=0.96$, accuracy plateaus at approximately 94.06%, corresponding to the performance of the single cheapest model (‘llama-3.1-8b’).
Overall, these results demonstrate CaMVo’s ability to balance exploration and exploitation, swiftly discover cost‐effective subsets, and reliably satisfy the target accuracy requirements.
To assess CaMVo’s sensitivity to dataset ordering, Figure 7 plots the mean cumulative accuracy and cost curves (solid lines) for $\delta=0.995$, $k_{\min}=1$, averaged over 20 random shuffles. Shaded regions indicate one standard deviation. Although the accuracy band starts wide, which reflects both the initial exploration and also the varying mixes of ’easier’ and ’harder’ instances; this rapidly contracts over time, confirming CaMVo’s reliable attainment of the target accuracy across permutations. The cost band likewise narrows, demonstrating stable convergence to low‐cost ensembles. Notably, the accuracy variability remains much smaller than the cost variability, as CaMVo does not optimize toward a fixed cost.
Figure 8 further investigates ordering effects by overlaying nine individual runs from these permutations. In every run, CaMVo exceeds the 95.14% accuracy target while driving per‐round cost below \$2.10, corroborating that its exploration–exploitation strategy, and the resulting cost–accuracy performance is effectively invariant to the input sequence.
Appendix F Experiments on the AG News Classification Dataset
We further evaluate CaMVo on the AG News Classification Dataset zhang2015character, which contains news articles gathered from over 2,000 online sources. Each article is categorized into one of four topics: World, Sports, Business, or Science/Technology. From the original training set of $120,000$ samples, we uniformly sample $50,000$ instances for our experiments. As before, we compare CaMVo against each individual LLM, a full‐ensemble Majority Vote, and the Baseline Method (Algorithm 2).
Models and setup.
We employ Anthropic’s Claude 3-7 Sonnet, and Claude 3-5 Haiku claude3sonnet; OpenAI’s o3-mini, o1-mini, and GPT-4o-mini openai2024models; and Meta’s LLaMA-3.3 and LLaMA-3.1 llama2024models. All queries use temperature = 0.3 and top-$p=1$, where applicable. We extract 384-dimensional contextual embeddings with all-MiniLM-L6-v2 wang2020minilm and approximate the confidence bound $\delta_{\mathcal{A}}(\bm{L},\bm{\omega})$ via the Beta‐CDF, as in §4.1. To improve computational efficiency, we again approximate the confidence score $\delta_{\mathcal{A}}(\bm{L},\bm{\omega})$ using the cumulative distribution function (CDF) of the Beta distribution rather than the closed-form expression in Lemma 2.1:
LLMs are queried using a consistent prompt format tailored for categorical output. The system instruction specifies the possible categories that the news article can be classified into, and the expected output format and behavior, to ensure that the model returns a single category. To form each instance for the User prompt of the LLM, we concatenate the article’s title and description provided in the dataset. The standard query format is shown below:
Query Format for AG News Classification Dataset System: Classify the following news article as WORLD, SPORT, BUSINESS, or SCIENCE/TECHNOLOGY. Respond with only the chosen category. If the article is ambiguous, select the closest matching category. User: <Title> : <Description>
An example query using this format, with a sample review from the AG News Classification Dataset, is provided below:
Example Query for AG News Classification Dataset System: Classify the following news article as WORLD, SPORT, BUSINESS, or SCIENCE/TECHNOLOGY. Respond with only the chosen category. If the article is ambiguous, select the closest matching category. User: Celtic beat Dunfermline 2-0: Celtic regained the top spot in the Scottish Premier League after a 2-0 victory away to Dunfermline. The result leaves Celtic top-of-the-table with 44 points from 18 games, a single point clear of arch-rivals Glasgow Rangers. Sentiment: LLM: SPORT
We apply a single random permutation to the $50,000$ sampled data instances from the dataset and maintain this identical ordering across all methods to ensure a fair and consistent comparison.
Results.
Table 8 (Right) reports the accuracy and cost (in dollars per million input tokens) of each LLM and the two baselines. Once again, the baseline underperforms the best individual model (85.68% vs. 87.16%) despite incurring a substantially higher cost. This performance gap may stem from the inherent ambiguity of the dataset, as certain news articles may belong to multiple categories, making consensus through majority voting more difficult to achieve.
Table 9 presents CaMVo’s accuracy–cost trade‐off across various thresholds $\delta$ and $k_{\min}\in\{1,3\}$. CaMVo’s hyperparameters are $\alpha=0.7$, $\lambda_{R}=5$, and $\lambda_{L}=1$; and the Target Accuracy is computed similarly as $\delta\times 85.68\%$. Across all configurations, CaMVo meets or exceeds its target accuracy. Further, CaMVo achieves less than half the cost (when $\delta=0.99$ and $k_{\min}=1$) at a slightly lower accuracy of $86.18\%$ compared to the baseline, confirming its practicality for large‐scale sentiment annotation without any pre‐training or ground-truth labels.
LLM / Method Accuracy (%) Cost
o3-mini 87.16 1.10
llama-3.3-70b 86.35 0.59
o1-mini 85.77 1.10
claude-3-7-sonnet 85.27 3.00
claude-3-5-haiku 84.71 0.80
gpt-4o-mini 82.63 0.15
llama-3.1-8b 77.85 0.05
Majority Vote 85.55 6.79
Baseline Method 85.68 6.79
Table 8: Accuracy and cost of individual LLMs and baseline ensemble methods on the AG News Classification Dataset.
CaMVo $\delta$
Target
Acc. (%)
Acc. (%)
$k_{\min}=1$
Cost
$k_{\min}=1$
Acc. (%)
$k_{\min}=3$
Cost
$k_{\min}=3$
0.999 85.59 85.98 6.82 85.98 6.82
0.998 85.51 85.98 6.82 85.98 6.82
0.995 85.25 86.36 5.24 86.36 5.24
0.99 84.82 86.18 2.98 86.18 2.98
0.98 83.97 84.72 1.39 84.72 1.39
0.96 82.25 83.38 0.52 83.85 0.99
0.95 81.40 83.18 0.40 83.77 0.96
0.9 77.11 80.66 0.29 83.75 0.88
0.85 72.83 79.62 0.26 83.74 0.87
0.80 68.54 79.34 0.23 83.73 0.87
Table 9: Accuracy and cost of CaMVo on the AG News Classification Dataset under varying confidence thresholds $\delta$ and $k_{\min}\in\{1,3\}$. For reference, the cost of the baseline method is \$6.79 per million tokens.
Figure 9: Cost–accuracy trade‐off for AG News Classification Dataset: gray dots show every LLM subset via weighted majority voting, yellow dots trace their Pareto‐optimal frontier, blue markers are CaMVo at $k_{\min}=1$, green markers at $k_{\min}=3$, cyan markers denote the individual single LLMs, and the red marker denotes the Baseline Method.
Figure 9 presents the analogous comparison of Figure 1 (Left) on the AG News Classification Dataset. As before, gray points and the yellow Pareto-frontier points show all possible subset combinations, cyan markers show individual single LLMs, while blue and green markers plot CaMVo at $k_{\min}=1$ and $3$, respectively. The red marker denotes the Baseline Method. Again, CaMVo closely matches the Pareto front in the low-cost regime (cost <0.8<0.8), but lags behind in higher-cost regions.